# From negative to positive $\Lambda$ through cosmological decreasing temperatures and its connection to spacetime foliation and string theory

### v1

@@ Abstract
String theories naturally predict a negative, while observations on the exponential expansion of the present Universe requires a positive value for the cosmological constant $\Lambda$. Solution to resolve this discrepancy is known in the framework of string theory and given for Asymptotically Safe (AS) quantum gravity by using Renormalization Group (RG). However, the former might describes unstable worlds and the latter lacks the inclusion of temperature which is otherwise crucial in the early Universe. Here we present a proposal for resolving this conflict by using a modified thermal RG method where the temperature parameter $T$ is given by the inverse radius of the compactified time-like dimension, similarly to spacetime foliation. In our scenario not the dimensionful temperature $T$, but the dimensionless temperature $\tau=T/k$ is kept constant when the RG scale $k$ is sent to zero and string theory is assumed to take place at very high while AS quantum gravity at intermediate and low temperatures. We show that the modified thermal RG study of AS quantum gravity models at very high temperatures results in a negative cosmological constant while turns it into a positive parameter for low temperatures.
@@ 1 Introduction
is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale, so it is not considered as the temperature, but rather understood as a running cutoff for thermal fluctuation. The dimensionless $\tau$ is identified as the physical temperature of the surrounding plasma. In this case the RG flow equations have real (and not pseudo) fixed points, which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations. In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points. In our fixed $\tau$ approach the dimensionful temperature is not defined, but the fixed points are. This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [12] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme. Moreover, in the recent work [11] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ aproach. The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [13] (i.e., non-Gaussian UV) fixed point. We showed that in the high temperature limit ($\tau\to\infty$) the $g$-coordinate of the Reuter fixed point vanishes and the cosmological constant takes on a negative value in the limit $k\to 0$. It is not in disagreement with observations, since during the thermal evolution of the Universe a thermal phase transition occurs and the cosmological constant runs to the expected positive value at low temperatures. This mechanism to solve the sign-problem of the cosmological constant is represented on Fig. 1.
@@
Motivated by the results in [11], here we study the generality of these findings. We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [14, 15]) has the cosmological constant with a negative (positive, repsectively) sign for large (small) temperature. Therefore, the scheme represented in Fig. 1 can be used in every case which can be seen as a natural solution for the sign problem of the cosmological constant. In particular, we study three extensions/variants of the simplest AS quantum gravity: the conformally reduced and the ghost-improved versions and its extension by scalar matter fields. In addition to that we discuss the connection to spacetime foliation.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged. The creation of this scientific communication was supported by the University of Miskolc with funding granted to the author I. Nándori within the framework of the institution’s Scientific Excellence Support Program. (Project identifier: ME-TKTP-2025-3).

### v2

# From negative to positive cosmological constant through decreasing temperature of the Universe: connection with string theory and spacetime foliation results

@@ Abstract
[…]
@@ 1 Introduction
[…]
Additional attempts achieving the anti-de Sitter (AdS) – de Sitter (dS) transition include considering two interacting dark energy fluids [5], taking running Barrow entropy into account [6], or examining quintessence fields with a negative cosmological constant [7].
@@
is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale, so it is not considered as the temperature, but rather understood as a running cutoff for thermal fluctuation. The dimensionless $\tau$ is identified as the physical temperature of the surrounding plasma. In this case the RG flow equations have real (and not pseudo) fixed points, which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations. In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points. In our fixed $\tau$ approach the dimensionful temperature is not defined, but the fixed points are. This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [17] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme. Moreover, in the recent work [16] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach. The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [18] (i.e., non-Gaussian UV) fixed point. We showed that in the high temperature limit ($\tau\to\infty$) the $g$-coordinate of the Reuter fixed point vanishes and the cosmological constant takes on a negative value in the limit $k\to 0$. It is not in disagreement with observations, since during the thermal evolution of the Universe a thermal phase transition occurs and the cosmological constant runs to the expected positive value at low temperatures. This mechanism to solve the sign-problem of the cosmological constant is represented on Fig. 1.
@@
Motivated by the results in [16], here we study the generality of these findings. We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [19, 20]) has the cosmological constant with a negative (positive, respectively) sign for large (small) temperature. Therefore, the scheme represented in Fig. 1 can be used in every case which can be seen as a natural solution for the sign problem of the cosmological constant. In particular, we study three extensions/variants of the simplest AS quantum gravity: the conformally reduced and the ghost-improved versions and its extension by scalar matter fields. In addition to that we discuss the connection to spacetime foliation.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged. The creation of this scientific communication was supported by the University of Miskolc with funding granted to the author I. Nándori within the framework of the institution’s Scientific Excellence Support Program. (Project identifier: ME-TKTP-2025-3). This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-376) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund.